# Advancing Spatio-Temporal Processing in Spiking Neural Networks through Adaptation

### v1

@@ 2.3 Stability analysis of discretized adLIF models
The upper frequency bound of adLIF neurons using the Euler-Forward discretization is illustrated in Fig. 3d. We can observe that for the Euler-Forward method, the maximum admissible frequency for stable dynamics converges toward zero as $\tau_{u}$ and $\tau_{w}$ increase (see Sections 4.7 in Methods for proof).
The difference in the stability properties has direct implications on network training: Firstly, since the neuron parameters $\tau_{u}$, $\tau_{w}$, $a$, and $b$ are usually trained, parameter ranges for Euler-Forward discretized adLIF neurons need to be constrained carefully to avoid stability issues. A similar issue has been observed in higuchi2024balanced and deckers2024co. Secondly, since SE discretization allows adLIF neurons to obtain oscillations of the full frequency bandwidth, they are theoretically capable of tuning to all possible frequencies present in the input.
2.4 Improved performance of SE-discretized adaptive RSNNs
@@
In the previous sections, we discussed the theoretical advantages of SE discretization over the more commonly used Euler-Forward discretization of the adLIF neuron model. From these theoretical considerations however, it is not directly evident whether this method provides any advantages in a realistic task setup. We therefore evaluated how recurrent networks of adLIF neurons compete against classical vanilla LIF networks as well as networks of adaptive neurons that have previously been considered in the literature: a model with threshold adaptation (ALIF) yinAccurateEfficientTimedomain2021, a constrained variant of the adLIF neuron model, similar to the model in this study, but with differences in discretization and neuron formulation (cAdLIF) deckers2024co, another adLIF network but with batch normalization (RadLIF) bittar2022surrogate, a feed-forward model with delays implemented as temporal convolutions (DCLS-Delays) hammouamri2023learning, and the balanced resonate-and-fire neuron model (BHRF) higuchi2024balanced, a variant of the resonate-and-fire neuron izhikevich2001resonate where output spikes do not disrupt the phase of the membrane potential oscillation. To that end, we constructed a recurrently connected SNN composed of one or two layers (depending on the task) of adLIF neurons, followed by a layer of leaky integrator (LI) neurons to provide a real-valued network output. We trained the SNN using BPTT with surrogate gradients leeTrainingDeepSpiking2016; bellec2018long; neftci2019surrogate; slayer. We used a dropout rate of $15\%$ except noted otherwise, but otherwise no regularization, normalization, or data augmentation methods. The trained parameters included the synaptic weights $\theta$, as well as all neuron parameters $a$, $b$, $\tau_{u}$ and $\tau_{w}$, which were not shared across neurons such that each neuron could have individual parameter values. Neurons were initialized heterogeneously, such that for each neuron the initial values of these parameters were chosen randomly from a uniform distribution over a pre-defined range. Heterogeneity has previously been shown to improve the performance of SNNs perez-nievesNeuralHeterogeneityPromotes2021. We applied a reparametrization technique for the training of time constants $\tau_{u}$ and $\tau_{w}$ and parameters $a$ and $b$, see Methods for details.
@@
We evaluated the networks on two commonly used audio benchmark datasets: Spiking Heidelberg Digits (SHD) cramerHeidelbergSpikingData2022 and Spiking Speech Commands (SSC) cramerHeidelbergSpikingData2022 , as well as an ECG dataset previously used to test ALIF neurons yinAccurateEfficientTimedomain2021. In Table 1 we report the test accuracy on the corresponding test sets. SHD does not define a dedicated validation set and previous work reported performances for networks validated on the test set, which is methodologically not clean. We therefore report results for two validation variants for SHD: with validation on the test set (to ensure comparability) and with validation on a fraction of the training set.
@@
SHD LIF $0.45$M $10$ $90.27\pm 0.73\%$
ALIF yinAccurateEfficientTimedomain2021 $0.14$M $1$ $90.4\%$
cAdLIF deckers2024co $38.7$k $10$ $94.19\%$
RadLIF bittar2022surrogate $3.9$M - $94.62\%$
DCLS-Delays hammouamri2023learning $0.2$M $10$ $95.07\pm 0.24\%$
EF-adLIF (2 layers) $0.45$M $20$ $94.68\pm 0.57\%$
SE-adLIF (1 layer) $37.5$k 20 $94.59\pm 0.27\%$
SE-adLIF (2 layers) $0.45$M 20 $95.81\pm 0.56\%$
SHD* BHRF higuchi2024balanced $0.1$M $5$ $92.7\pm 0.7\%$
SE-adLIF (1 layer) $37.5$k 20 $93.18\pm 0.74\%$
SE-adLIF (2 layers) $0.45$M 20 $93.79\pm 0.76\%$
SSC LIF $1.7$M $1$ $75.23\%$
ALIF yinAccurateEfficientTimedomain2021 $0.78$M $1$ $74.2\%$
RadLIF bittar2022surrogate $3.9$M - $77.4\%$
cAdLIF deckers2024co $0.35$M $1$ $77.5\%$
DCLS-Delays hammouamri2023learning $2.5$M $5$ $80.69\pm 0.21\%$
EF-adLIF (2 layers) $1.6$M $5$ $79.80\pm 0.18\%$
SE-adLIF (2 layers) $1.6$M $5$ $80.44\pm 0.26\%$
ECG LIF (1 layer) $1.8$k 5 $77.8\pm 1.85\%$
ALIF yinAccurateEfficientTimedomain2021 $1.8$k $1$ $85.9\%$
BHRF higuchi2024balanced $1.8$k $5$ $87.0\pm 0.4\%$
EF-adLIF (1 layer) $1.8$k $20$ $87.10\pm 0.46\%$
SE-adLIF (1 layer) $1.8$k $20$ $86.88\pm 0.40\%$
EF-adLIF (2 layers) $4.6$k $20$ $87.83\pm 0.51\%$
SE-adLIF (2 layers) $4.6$k $20$ $88.18\pm 0.36\%$
@@
Our results show that recurrent adLIF networks clearly outperformed LIF network baselines on all tasks. For all considered datasets, recurrent SE-discretized adLIF networks performed better than previously considered recurrent SNNs. For SSC, their performance was slightly below that of the DCLS-model hammouamri2023learning, a feed-forward network using extensive delays trained via dilated convolutions. Networks composed of SE-discretized adLIF neurons (SE-adLIF networks) performed significantly better than those based on Euler-forward discretization (EF-adLIF networks) on SHD and SSC (significance values for a two-tailed t-test were $p<0.000001$ for SHD and $p<0.005$ for SSC). Small networks with a single recurrent layer on ECG performed on-par ($p=0.115$), while SE-adLIF networks significantly improved over EF-adLIF networks when larger networks with two recurrent layers were used ($p<0.02$). We found that EF-adLIF networks suffered from severe instabilities if neuron parameters were not constrained to values in which the decay rate exceeds the critical boundary of $r=1$, resulting in instabilities for example in the ECG task, see Section A in Appendix. The SE method is hence the preferred choice when adLIF neurons are used in a discretized form.
@@
AdLIF neurons could, depending on their parameters, exhibit many different experimentally observed neuronal dynamics gerstner2014neuronal. We wondered whether networks trained on spatio-temporal classification tasks utilized the diverse dynamical behaviors of adLIF neurons. To that end, we investigated the resulting parameterizations of adLIF neurons in networks trained on SHD, and indeed found a heterogeneous landscape of neuron parameterizations, see Section B in the Appendix.
2.5 AdLIF networks accurately predict trajectories of dynamical systems
@@
2.6 Computational properties of adLIF networks
@@
Figure 6: Temporal feature detection in adLIF networks. a) Two samples of classes $2$ and $17$ of the burst sequence detection (BSD) task (see main text). b) Classification error of adLIF and LIF networks with equal parameter count for different numbers of classes in the BSD task. c) Schematic illustration of network feature visualization. An initial noise sample $\bm{x_{0}}$ is passed through a trained network with frozen network parameters. The classification loss of the network output with respect to some predefined target class $c$ is computed and back-propagated through the network to obtain the gradient $\nabla_{X}L(X,c)|_{X^{0}}$ of the loss with respect to input $X^{0}$. This gradient is applied to the sample and the procedure is repeated to obtain a final sample $X^{K}$ after $K=400$ iterations. d) Samples generated by the feature visualization procedure from panel c from networks trained on the $20$-class BSD task. White dots denote the locations of the class-descriptive bursts. We generated samples for classes $2$ and $17$ which were the most misclassified classes of LIF and adLIF networks respectively. e) Samples generated from an adLIF network and a LIF network trained on SHD for different target classes $c$ (top) and the corresponding network output over time (bottom). The grey shaded area (at $t>100$) denotes the time span relevant for the loss, all outputs before this time span were ignored, see Methods for details.
This task is motivated from neuroscientific experiments which show the importance of sequences of transient increases of spike rates in cortex harvey2012choice; LISMAN199738. A class in this task is defined by a specific pre-defined temporal sequence of bursts across a fixed sub-population of three of these neurons. Other neurons emit a random burst at a random time each and additionally, neurons fire with a background rate of $50$ Hz. Bursts were implemented as smooth, transient increases of firing rate resulting in approximately $7$ spikes per burst, see Section 4.12 in Methods for details. We tested single-layer recurrent adLIF networks ($510$ neurons) and single-layer recurrent LIF networks with the same number of trainable parameters. AdLIF networks clearly outperformed LIF networks on this task, see Fig. 6b. For the case of $10$ classes, the adLIF network reached an average classification error of $2.31\%$ on the test set, whereas the LIF network only achieved a test error of $6.96\%$ despite a low training error ($<2\%$).
@@
where $\eta$ is the update step size, $\nabla_{X}f(X)|_{X^{k}}$ denotes the gradient of $f$ with respect to $X$ evaluated at $X^{k}$, and $\zeta_{k}$ is a normalization factor, see Fig. 6c and Section 4.13 in Methods for details. After each update, we applied additional regularization to the data sample $X^{k+1}$ (see Methods for details). We repeated this procedure for $K=400$ iterations, such that the final $X^{K}$ yielded a very strong prediction for class $c$.
We performed this feature visualization for the BSD task and for the SHD task. Fig. 6d shows the resulting artificial samples $X^{K}$ for an adLIF network and a LIF network trained on the BSD task. The class-defining burst timings are indicated as black circles with white filling.
The adLIF-generated samples exhibited a strong temporal structure that captures the relevant temporal structure of the class. This can be observed visually by comparing the position of strong activations with the class-defining burst timings. In contrast, the samples generated from the trained LIF network displayed less precise resemblance of class-descriptive features, and showed less temporal variation. This gap of specificity of the features in the generated samples of adLIF versus LIF networks might explain the performance gap between these: While the less precise temporal tuning of LIF networks suffices to achieve a high accuracy on the training data, it falls behind in terms of generalization on the test set, due to confusion of temporal features from different classes. Interestingly, the same analysis for an ALIF network yinAccurateEfficientTimedomain2021 with threshold adaptation revealed that the temporal tuning of ALIF networks is comparable to that of LIF networks, indicating the importance of oscillatory dynamics for temporal feature detection, see Section D in the Appendix.

### v3

@@ Results
The upper frequency bound of adLIF neurons using the Euler-Forward discretization is illustrated in Fig. 3f. We can observe that for the Euler-Forward method, the maximum admissible frequency for stable dynamics converges toward zero as $\tau_{u}$ and $\tau_{w}$ increase (see Section ”Stable ranges for intrinsic frequencies of EF-adLIF” in ”Methods” for proof).
An immediate advantage of using the SE-adLIF over EF-adLIF is the guaranteed stability over the entire range of possible oscillation frequencies. This property in particular comes into play when the neuron parameters $\tau_{u}$, $\tau_{w}$, $a$, and $b$ are trained. While for most tasks only a sub-range of this viable frequency range might be required, it is guaranteed that the SE-adLIF neuron is stable for any such frequency. In other words, for SE-adLIF neurons with frequencies below the Nyquist frequency, no unstable parameter configurations exists. This is not the case for the EF-adLIF neuron: even in instances of very low oscillation frequencies the stability boundary might be overshot (see Fig. 3c,d). Later in the manuscript (see Section ”Accurate prediction of dynamical system trajectories” and Fig. 4g), we discuss the relationship between the neuron frequency range and the performance in an oscillator regression task. In our simulations, we clip $a$ to fixed constant upper and lower bounds, given by task-dependent hyperparameters, independent of parameters $\tau_{u}$ and $\tau_{w}$. While this constraint suffices for most tasks, it introduces a trade-off between the decay of the neuron and the oscillation frequency. This can be observed in Fig. 3c (right), where neurons with a high frequency are restricted to a fast decay. A trivial extension to the SE-adLIF model would be to dynamically adjust the upper bound for parameter $a$ with respect to parameters $\tau_{u}$ and $\tau_{w}$ by computing $a_{\mathrm{max}}^{\text{SE}}$ (as defined by Eq. (68) in ”Methods”) after each training step and clipping $a$ to the interval $[0,a_{\mathrm{max}}^{\text{SE}}]$. This would allow the neuron model to utilize the entire frequency range for any combination of $\tau_{u}$ and $\tau_{w}$. This extension is not possible for the EF-adLIF neuron, since using $a_{\mathrm{max}}^{\text{EF}}$ (as defined by Eq. (34) in ”Methods”) as an upper bound instead of a constant value would still not allow the neuron to use the entire frequency range. This exact case, where $a$ is set to $a_{\mathrm{max}}^{\text{EF}}$, is shown in Fig. 3f.
The favorable stability properties induced by the SE discretization should generalize well to other neuron models with two bi-directionally coupled neuron states. Two examples for such neuron models are the adaptive exponential integrate-and-fire (AdEx) bretteAdaptiveExponentialIntegrateandFire2005 model and the Balanced Harmonic Resonate-and-Fire (BHRF) higuchi2024balanced model. For the latter, we observed that applying the SE-discretization not only alleviates the necessity of the frequency-dependent divergence boundary, which was introduced by the authors to ensure stability of the model, but also recovers the direct relationship between neuronal parameters $\omega$ and $b$ and the effective oscillation frequency $\omega_{\text{eff}}$ and effective damping coefficient $b_{\text{eff}}$ of the discretized neuron. Details can be found in Supplementary Notes 1 and Supplementary Figure 1.
Improved performance of SE-discretized adaptive RSNNs
We first evaluated how recurrent networks of the described adLIF neurons perform in comparison to classical vanilla LIF networks. We compared LIF and SE-adLIF networks on two commonly used audio benchmark datasets: Spiking Heidelberg Digits (SHD) cramerHeidelbergSpikingData2022 and Spiking Speech Commands (SSC) cramerHeidelbergSpikingData2022 , as well as an ECG dataset previously used to test ALIF neurons yinAccurateEfficientTimedomain2021. For LIF baselines, we used results from previously reported studies as well as additional simulations to ensure identical setups and comparable parameter counts.
@@
We obtained our results by constructing a recurrently connected SNN composed of one or two layers (depending on the task) of adLIF (resp. LIF) neurons, followed by a layer of leaky integrator (LI) neurons to provide a real-valued network output. We trained both the adLIF and LIF SNNs using BPTT with surrogate gradients leeTrainingDeepSpiking2016; bellec2018long; neftci2019surrogate; slayer. We used a dropout rate of $15\%$ except noted otherwise, but otherwise no regularization, normalization, or data augmentation methods. The trained parameters included the synaptic weights $\theta$, as well as all neuron parameters $a$, $b$, $\tau_{u}$ and $\tau_{w}$ in the SE-adLIF case, and membrane time constants $\tau$ in the LIF case. The parameters were not shared across neurons such that each neuron could have individual parameter values. Neurons were initialized heterogeneously, such that for each neuron the initial values of these parameters were chosen randomly from a uniform distribution over a pre-defined range. Heterogeneity has previously been shown to improve the performance of SNNs perez-nievesNeuralHeterogeneityPromotes2021. We applied a reparametrization technique for the training of time constants $\tau_{u}$ and $\tau_{w}$ and parameters $a$ and $b$ for the SE-adLIF model, as well as the membrane time constants $\tau$ for the LIF models, see ”Methods” for details.
Table 1 summarizes both the baselines from prior studies and our own results. While LIF networks in our simulations performed better than all previously reported LIF baselines, they still performed significantly worse than the SE-adLIF networks with the same or lesser parameter counts across all tasks. This result provides clear empirical support for the superiority of adLIF networks over LIF networks, both in terms of parameter efficiency and overall performance.
Table 1: Comparison of recurrent LIF and adLIF networks on spike-encoded speech recognition datasets. The ’N.A.’ indicates lack of information from the authors. ’This work’ indicates our experimental results. The reported results correspond to the accuracy on the corresponding test set.
Model Publication #Params Test Acc. [%]
SHD LIF Cramer et al. 2022 cramerHeidelbergSpikingData2022 N.A. $83.2\pm 1.3$
LIF Deckers et al. 2024 deckers2024co $37.9$k $84.49$
LIF Bittar & Garner 2022 bittar2022surrogate $141$k $87.04$
LIF Bittar & Garner 2022 bittar2022surrogate $3.8$M $89.29$
LIF This work $450$k $90.27\pm 0.73$
SE-adLIF This work $450$k $\mathbf{95.81\pm 0.56}$
SSC LIF Cramer et al. 2022 cramerHeidelbergSpikingData2022 N.A. $50.9\pm 1.1$
LIF Deckers et al. 2024 deckers2024co $0.34$M $71.76$
LIF Bittar & Garner 2022 bittar2022surrogate $141$k $66.67$
LIF Bittar & Garner 2022 bittar2022surrogate $1.1$M $68.14$
LIF This work $1.7$M $75.23$
SE-adLIF This work $1.6$M $\mathbf{80.44\pm 0.26}$
ECG LIF This work $1.8$k $77.8\pm 1.85$
SE-adLIF This work $1.8$k $\mathbf{86.88\pm 0.40}$
In the previous sections, we discussed the theoretical advantages of SE discretization over the more commonly used Euler-Forward discretization of the adLIF neuron model. The discussed theoretical advantages of SE, for example near-independence of the neuron dynamics from the discretization time step $\Delta t$ or the closer adherence to the continuous model, give indirect practical benefits when dealing with such neurons. It is not clear however, whether the utilization of the SE discretization can provide an improvement in performance over the commonly used EF method. To answer this question, we not only compared both the EF-adLIF and the SE-adLIF model to each other, but also to state-of-the-art spiking neural networks: a model with threshold adaptation (ALIF) yinAccurateEfficientTimedomain2021, a constrained variant of the adLIF neuron model, similar to the model in this study, but with differences in discretization and neuron formulation (cAdLIF) deckers2024co, another adLIF network but with batch normalization (RadLIF) bittar2022surrogate, a feed-forward model with delays implemented as temporal convolutions (DCLS-Delays) hammouamri2023learning, and the balanced resonate-and-fire neuron model (BHRF) higuchi2024balanced, a variant of the resonate-and-fire neuron izhikevich2001resonate where output spikes do not disrupt the phase of the membrane potential oscillation.
@@
In Table 2 we report the test accuracy of the various models on the corresponding test sets. SHD does not define a dedicated validation set and previous work reported performances for networks validated on the test set, which is methodologically not clean. We therefore report results for two validation variants for SHD: with validation on the test set (to ensure comparability) and with validation on a fraction of the training set.
@@
SHD LIF $0.45$M $10$ $90.27\pm 0.73$
ALIF yinAccurateEfficientTimedomain2021 $0.14$M $1$ $90.4$
cAdLIF deckers2024co $38.7$k $10$ $94.19$
RadLIF bittar2022surrogate $3.9$M - $94.62$
DCLS-Delays hammouamri2023learning $0.2$M $10$ $95.07\pm 0.24$
EF-adLIF (2 layers) $0.45$M $20$ $94.68\pm 0.57$
SE-adLIF (1 layer) $37.5$k 20 $94.59\pm 0.27$
SE-adLIF (2 layers) $0.45$M 20 $95.81\pm 0.56$
SHD* BHRF higuchi2024balanced $0.1$M $5$ $92.7\pm 0.7$
SE-adLIF (1 layer) $37.5$k 20 $93.18\pm 0.74$
SE-adLIF (2 layers) $0.45$M 20 $93.79\pm 0.76$
SSC LIF $1.7$M $1$ $75.23$
ALIF yinAccurateEfficientTimedomain2021 $0.78$M $1$ $74.2$
RadLIF bittar2022surrogate $3.9$M - $77.4$
cAdLIF deckers2024co $0.35$M $1$ $77.5$
DCLS-Delays hammouamri2023learning $2.5$M $5$ $80.69\pm 0.21$
EF-adLIF (2 layers) $1.6$M $5$ $79.80\pm 0.18$
SE-adLIF (2 layers) $1.6$M $5$ $80.44\pm 0.26$
ECG LIF (1 layer) $1.8$k 5 $77.8\pm 1.85$
ALIF yinAccurateEfficientTimedomain2021 $1.8$k $1$ $85.9$
BHRF higuchi2024balanced $1.8$k $5$ $87.0\pm 0.4$
EF-adLIF (1 layer) $1.8$k $20$ $87.10\pm 0.46$
SE-adLIF (1 layer) $1.8$k $20$ $86.88\pm 0.40$
EF-adLIF (2 layers) $4.6$k $20$ $87.83\pm 0.51$
SE-adLIF (2 layers) $4.6$k $20$ $88.18\pm 0.36$
@@
For all considered datasets, recurrent SE-discretized adLIF networks performed better than previously considered recurrent SNNs. For SSC, their performance was slightly below that of the DCLS model hammouamri2023learning, a feed-forward network using extensive delays trained via dilated convolutions. Unlike our model, however, DCLS employs temporal convolutions to implement delays and incorporates batch normalization. These properties make the DCLS model less suitable for neuromorphic use cases. Nevertheless, we included it in our results table for comparison, as the delays in neural connectivity provide an interesting orthogonal complement to the enhanced somatic dynamics of the models studied in our work (see also ”Discussion”). Networks composed of SE-discretized adLIF neurons (SE-adLIF networks) performed significantly better than those based on Euler-forward discretization (EF-adLIF networks) on SHD and SSC (significance values for a two-tailed t-test were $p<0.000001$ for SHD and $p<0.005$ for SSC). Small networks with a single recurrent layer on ECG performed on-par ($p=0.115$), while SE-adLIF networks significantly improved over EF-adLIF networks when larger networks with two recurrent layers were used ($p<0.02$). We found that EF-adLIF networks suffered from severe instabilities if neuron parameters were not constrained to values in which the decay rate exceeds the critical boundary of $r=1$, resulting in instabilities for example in the ECG task, see Supplementary Notes 2 and Supplementary Table 1. The SE method is hence the preferred choice when adLIF neurons are used in a discretized form.
@@
AdLIF neurons could, depending on their parameters, exhibit many different experimentally observed neuronal dynamics gerstner2014neuronal. We wondered whether networks trained on spatio-temporal classification tasks utilized the diverse dynamical behaviors of adLIF neurons. To that end, we investigated the resulting parameterizations of adLIF neurons in networks trained on SHD, and indeed found a heterogeneous landscape of neuron parameterizations, see Supplementary Notes 3 and Supplementary Figure 2.
Accurate prediction of dynamical system trajectories
@@
High-fidelity neuromorphic audio compression
In all experiments so far we observed a significant superiority of the adLIF neuron over the LIF neuron, both in terms of parameter efficiency and overall performance. Yet, it is unclear how these observed improvements transfer from benchmarks and toy tasks to real-world neuromorphic applications. To take a step towards answering this question, we compared the performance of adLIF and LIF neurons in the task of raw audio compression. The goal of this task is to first compress and then transmit a raw audio signal as energy-efficient as possible, while sacrificing as little signal quality as possible. Similar setups with neuromorphic processing and spike-based transmission have been discussed in various studies rothSpikeBasedSensingCommunication2022; wuNeuromorphicWirelessSplit2024; ke2024; chenNeuromorphicWirelessCognition2023a; savazziMutualInformationAnalysis2024 as promising research direction for low-power IoT applications with smart wireless sensors. Our study is, to the best of our knowledge, the first to address audio compression using plain RSNNs that deliberately exclude batch normalization, temporal convolutions, and transformer-based architectures to maintain compatibility with standard neuromorphic processors.
Figure 5: Audio compression task setup and results. a) (top) The conceptual hardware setup considered in the audio compression task. See main text and ”Methods”  for details. (bottom) Architecture of our simulated SNNs, closely resembling the conceptual neuromorphic setup. The raw waveform input to the encoder RSNN, the $16$-dimensional spike train between encoder and decoder, as well as the reconstructed waveform were taken from our simulations with SE-adLIF networks. b) Comparison between the ground truth and the reconstructed waveform for the LIF (top), EF-adLIF (middle), and SE-adLIF network (bottom).
The setup is illustrated in Fig. 5a: We conceptually consider a small device containing a digital neuromorphic chip that receives raw, unprocessed audio from a microphone. For our study, we used SNN simulations and did not implement this setup in real hardware, but simulated the SNNs that would run on such chips, see Fig. 5a, bottom. In the conceptual setup, the chip processes the waveform by implementing a small SNN with a bottleneck output layer consisting of very few neurons. It sends the spike-encoded audio data through a sparse wireless communication channel savazziMutualInformationAnalysis2024 to a receiving device. This receiving device, a second neuromorphic processor, could in principle perform arbitrary post-processing on the spike-encoded data. In our simulations we considered the most general case, which is the reconstruction of the ground truth waveform from the sparse spikes. For the spike-based communication we assume low-latency pulse-driven radio transmission, for example IR-UWB fernandesRecentAdvancesIRUWB2010, that features adaptive energy consumption, depending on the presence of input signal. In the absence of an input signal (i.e. silence), almost no energy is consumed by the transmitting device, hence this technology is a promising candidate for ultra low-power neuromorphic sensing devices chenNeuromorphicWirelessCognition2023a. In contrast, if conventional frame-based digital transmission is used, the transmission rate is constant and power is consumed at a constant rate. In pulse-driven spike encoding, the timing of spikes is implicitly encoded by the timing of the emitted radio pulse, such that the spike timing does not need to be explicitly transmitted as payload.
Audio compression requires balancing the quality of the reconstructed signal against the data transmission rate at the bottleneck. In our simulations, we constrained the encoder SNN to very sparse spiking activity at the output layer to achieve low-bandwidth transmission. This aligns with our goal of ultra-low-energy processing, as energy consumption in neuromorphic systems is directly tied to spike rate. We considered as few as $16$ output neurons for the encoder, regularized to not exceed a total maximum spike rate of $6$k spikes per second. Assuming that a single spike is equivalent to $1$ bit, the upper regularization bound for the average data transmission rate between encoder and decoder is $6$ kbps. Under this constraint, we compared the quality of the reconstructed audio signal of LIF and adLIF neurons with common audio compression codecs valin2012definition; dietz2015overview and the state-of-the-art Residual Vector Quantization (RVQ) method defossez2022high evaluated at the same bandwidth of $6$ kbps. The results are shown in Table 3. Details on the task setup and the simulations can be found in Section ”Details for the audio compression task” in ”Methods”. We provide uncompressed audio samples (Supplementary Audio 1) and reconstructions using the LIF (Supplementary Audio 2), EF-adLIF (Supplementary Audio 3), and SE-adLIF networks (Supplementary Audio 4) as supplements to this article. Our findings indicate that adaptive LIF neurons achieve a significantly higher reconstruction quality than vanilla LIF neurons (see Fig. 5b for an example). Moreover, SE-discretized adLIF neurons yield significantly higher reconstruction quality, as demonstrated by both quantitative metrics (Table 3) and visual waveform comparisons (Fig. 5b). Given that waveform data spans a broad frequency spectrum, the stability of the SE-adLIF neurons over the entire frequency range — as discussed in Section ”Stability analysis of discretized adLIF models” — offers a clear advantage over EF-adLIF. Since LIF neurons lack intrinsic membrane potential oscillations, they depend heavily on recurrent network dynamics to detect, encode and decode oscillatory patterns in their input. We observed the same phenomenon in the oscillatory task in Fig. 4, where LIF neurons performed poorly. While the SE-adLIF model achieved the best performance in terms of scale-invariant signal-to-noise ratio luo2018tasnet, the state-of-the-art neural network model RVQ exhibits better performance on the VISQOL measure hines2015visqol, but at the cost of a much larger model with a >27×>27$\times$> 27 × increase in the number of parameters. The compactness of the SE-adLIF model allowed audio compression and decoding in 1.3×1.3$\times$1.3 × real-time on a consumer single tread CPU (AMD Ryzen 7 5800H). In Table 3, we also report performances of two standard audio codecs (OPUS and EVS) on our test set, showing that the SE-adLIF networks achieve competitive performance. In the next few sections, we further explore the inductive bias introduced by oscillatory membrane potentials.
Table 3: Performance comparison between LIF, EF-adLIF, and SE-adLIF networks with state-of-the-art audio codecs in the audio compression task. We used the same network layout for all neuron models and report mean and standard devivation on the LibriTTS clean test set zen2019libritts across $4$ runs except for LIF models, for which only 2 runs converged and were taken into account. See Section ”Details for the audio compression task” in ”Methods”  for details. SI-SNR denotes the scale-invariant signal-to-noise ratio luo2018tasnet, VISQOL the value obtained from the Virtual Speech Quality Objective Listener hines2015visqol. OPUS and EVS are audio codecs, hence no parameter counts are provided. Sam. Rate refers to the respective sampling rate of the audio signal. Avg. Bandw. refers to the average number of spikes per second (Sp./s) at the bottleneck layer for the spiking networks, and the data transmission rate for the other, frame-based methods.
Model Sam. Rate #Params Avg. Bandw. SI-SNR luo2018tasnet VISQOL hines2015visqol
[…]
Computational properties of adLIF networks
@@
Figure 7: Temporal feature detection in adLIF networks. a) Two samples of classes $2$ and $17$ of the burst sequence detection (BSD) task (see main text). b) Classification error of adLIF and LIF networks with equal parameter count for different numbers of classes in the BSD task. c) Schematic illustration of network feature visualization. An initial noise sample $\bm{x_{0}}$ is passed through a trained network with frozen network parameters. The classification loss of the network output with respect to some predefined target class $c$ is computed and back-propagated through the network to obtain the gradient $\nabla_{X}L(X,c)|_{X^{0}}$ of the loss with respect to input $X^{0}$. This gradient is applied to the sample and the procedure is repeated to obtain a final sample $X^{K}$ after $K=400$ iterations. d) Samples generated by the feature visualization procedure from panel c from networks trained on the $20$-class BSD task. White dots denote the locations of the class-descriptive bursts. We generated samples for classes $2$ and $17$ which were the most misclassified classes of LIF and adLIF networks respectively. e) Samples generated from an adLIF network and a LIF network trained on SHD for different target classes $c$ (top) and the corresponding network output over time (bottom). The grey shaded area (at $t>100$) denotes the time span relevant for the loss, all outputs before this time span were ignored, see ”Methods” for details.
This task is motivated from neuroscientific experiments which show the importance of sequences of transient increases of spike rates in cortex harvey2012choice; LISMAN199738. A class in this task is defined by a specific pre-defined temporal sequence of bursts across a fixed sub-population of three of these neurons. Other neurons emit a random burst at a random time each and additionally, neurons fire with a background rate of $50$ Hz. Bursts were implemented as smooth, transient increases of firing rate resulting in approximately $7$ spikes per burst, see Section ”Details for the Burst Sequence Detection (BSD) task” in ”Methods” for details. We tested single-layer recurrent adLIF networks ($510$ neurons) and single-layer recurrent LIF networks with the same number of trainable parameters. AdLIF networks clearly outperformed LIF networks on this task, see Fig. 7b. For the case of $10$ classes, the adLIF network reached an average classification error of $2.31\%$ on the test set, whereas the LIF network only achieved a test error of $6.96\%$ despite a low training error ($<2\%$).
@@
where $\eta$ is the update step size, $\nabla_{X}f(X)|_{X^{k}}$ denotes the gradient of $f$ with respect to $X$ evaluated at $X^{k}$, and $\zeta_{k}$ is a normalization factor, see Fig. 7c and Section ”Optimization-based feature visualization” in ”Methods” for details. After each update, we applied additional regularization to the data sample $X^{k+1}$ (see ”Methods” for details). We repeated this procedure for $K=400$ iterations, such that the final $X^{K}$ yielded a very strong prediction for class $c$. We performed this feature visualization for the BSD task and for the SHD task. Fig. 7d shows the resulting artificial samples $X^{K}$ for an adLIF network and a LIF network trained on the BSD task. The class-defining burst timings are indicated as black circles with white filling.
The adLIF-generated samples exhibited a strong temporal structure that captures the relevant temporal structure of the class. This can be observed visually by comparing the position of strong activations with the class-defining burst timings. In contrast, the samples generated from the trained LIF network displayed less precise resemblance of class-descriptive features, and showed less temporal variation. This gap of specificity of the features in the generated samples of adLIF versus LIF networks might explain the performance gap between these: While the less precise temporal tuning of LIF networks suffices to achieve a high accuracy on the training data, it falls behind in terms of generalization on the test set, due to confusion of temporal features from different classes. Interestingly, the same analysis for an ALIF network yinAccurateEfficientTimedomain2021 with threshold adaptation revealed that the temporal tuning of ALIF networks is comparable to that of LIF networks, indicating the importance of oscillatory dynamics for temporal feature detection, see Supplementary Notes 4 and Supplementary Figure 4.